OpenStudy (anonymous):

Dameon solved this problem. Adrian paid $465.60 for 120 rubber pads for the bottom of his store's chairs. How much did Adrian pay for each rubber pad? Dameon thought the answer was $3.88. Is Dameon's answer reasonable? A. No, Dameon's answer should be closer to $0.40 B. No, Dameon's answer should be closer to $40. C. No, Dameon's answer should be closer to $400. D. Yes, Dameon's answer is reasonable.

OpenStudy (danjs):

120 cost that much, so 1 will cost (1/120)the that cost

OpenStudy (danjs):

one hundred twentieth the cost

OpenStudy (anonymous):

??

OpenStudy (danjs):

\[\frac{ $X~}{ 1~pad}=\frac{$465.60 }{ 120~pads }\]

OpenStudy (anonymous):

is it C?

OpenStudy (danjs):

X cost for 1 pad is equal to 465.60 cost for 120 pads

OpenStudy (danjs):

Divide 465.60 by 120, that is the cost per 1 pad

OpenStudy (danjs):

\[\frac{ $465.60 }{ 120~pads } = \frac{ [$465.60/120] }{1~pad }\]

OpenStudy (danjs):

or $3.88 / 1 pad

OpenStudy (anonymous):

so it would be c?

OpenStudy (anonymous):

never mind

OpenStudy (danjs):

no, if he paid $400 for each one, then he paid 120 x 400 for the lot of em, which is not $465.60

OpenStudy (anonymous):

so what is it?

OpenStudy (danjs):

3.88 per , same as Dameon calculated